# CAML: Collaborative Auxiliary Modality Learning for Multi-Agent Systems

### v2

@@ 4.1 Collaborative Decision-Making
Teacher 96.0$\pm$0.3 86.8$\pm$0.4 73.3$\pm$0.8 81.2$\pm$0.6 68.8$\pm$0.9 85.8$\pm$0.7
RGB-only BC 85.2$\pm$1.0 83.2$\pm$0.9 56.4$\pm$1.5 78.0$\pm$1.2 55.8$\pm$1.3 81.0$\pm$1.2
CAML 92.8$\pm$0.8 85.5$\pm$0.5 67.3$\pm$1.0 79.7$\pm$1.1 66.1$\pm$1.4 83.0$\pm$1.3
@@ 4.2 Collaborative Semantic Segmentation
To further evaluate our approach, we focus on collaborative semantic segmentation by conducting experiments with real-world data from aerial-ground robots. We use the dataset CoPeD [41], with one aerial robot and one ground robot, in two different real-world scenarios of the indoor NYUARPL and the outdoor HOUSEA. For more details about the dataset, please refer to [41]. Additionally, we introduce noise to the RGBD data collected by the ground robot. For both aerial and ground robots, RGB and depth data are used during training, while only RGB data is used during testing in CAML.
@@
We adopt the FCN [23] architecture as the backbone for semantic segmentation. To process RGB and depth data locally for each robot, we use ResNet-18 [14] as the encoder to extract feature maps of size $7\times 7$. Please refer to Appendix A.3.1 for more details of the experimental setup. We first train a teacher model offline with aerial-ground robots collaboration using cross-entropy loss, where each robot has both RGB and depth data. Then we train a student model to mimic the behavior of the teacher model with only RGB data for both aerial and ground robots through KD. The KD process is similar to that of the collaborative decision-making in CAV, but here we use a cross-entropy loss as the student task loss. Please see Appendix A.4 for the detailed training settings. We evaluate performance using the Mean Intersection over Union (mIoU) metric, which quantifies the average overlap between predicted segmentation outputs and ground truth across all classes. We compare the performance of CAML with other baselines including AML [32] and FCN [23]. In the AML approach, only the ground robot operates, with RGB and depth data available during training but only RGB data used for testing. The FCN approach involves only the ground robot operating with RGB data for both training and testing.
Experimental Results.
We first present the experimental results of baselines comparison in Table 3, where CAML demonstrates superior performance in terms of mIoU across both indoor and outdoor environments. Specifically, CAML achieves an improvement of mIoU for ${\bf 8.9\%}$ in indoor scenario and ${\bf 10.6\%}$ in outdoor scenario compared to AML [32]. We also show the qualitative results in Fig. 5. As we can see, despite the noisy input image from the ground robot, CAML produces predictions that are closest to the ground truth. This improvement is attributed to CAML’s multi-agent collaboration, which provides complementary information to enhance data coverage and offers a more comprehensive understanding of the scenes. Additionally, the utilization of auxiliary depth data during training results in more precise segmentation outputs.
Table 3: Baseline Comparison of Semantic Segmentation on real-world dataset CoPeD [41] using aerial-ground robots in indoor and outdoor environments. CAML achieves the highest mIoU in both environments, with upto 10.6% higher accuracy.
@@ 5 Conclusions
In conclusion, we propose Collaborative Auxiliary Modality Learning (CAML), a unified framework for multi-modal multi-agent systems. Unlike prior methods that either focus on multi-agent collaboration without modality reduction or address multi-modal learning in single-agent settings, CAML integrates both aspects. It enables agents to collaborate using shared modalities during training while allowing efficient, modality-reduced inference. This not only lowers computational costs and data requirements at test time but also enhances predictive accuracy through multi-agent collaboration. We provide an intuitive analysis of CAML in terms of data coverage and information gain, justifying its effectiveness. CAML demonstrates up to a $58.1\%$ improvement in accident detection for connected autonomous driving in complex scenarios and up to a $10.6\%$ mIoU gain in real-world aerial-ground collaborative semantic segmentation. These improvements underscore the practical implications of our framework, especially for resource-constrained environments. For limitations and future work, please see Appendix A.6.
@@ A.2.1 Dataset Details
The dataset is generated using the AUTOCASTSIM [29] benchmark, which features three complex and accident-prone traffic scenarios for connected autonomous driving, characterized by limited sensor coverage or obstructed views. These scenarios are realistic and include background traffic of 30 vehicles. They involve challenging interactions such as overtaking, lane changing, and red-light violations, which inherently increase the risk of accidents: (1) Overtaking: A sedan is blocked by a truck on a narrow, two-way road with a dashed centerline. The truck also obscures the sedan’s view of oncoming traffic. The ego vehicle must decide when and how to safely pass the truck. (2) Left Turn: The ego vehicle attempts a left turn at a yield sign. Its view is partially blocked by a truck waiting in the opposite left-turn lane, reducing visibility of vehicles coming from the opposite direction. (3) Red Light Violation: As the ego vehicle crosses an intersection, another vehicle runs a red light. Due to nearby vehicles waiting to turn left, the ego vehicle’s sensors are unable to detect the violator.
Following the setup established by prior works COOPERNAUT [5] and STGN [8], we collect 24 data trails for each scenario, using 12 trails for training and the remaining 12 for testing. Each trail represents a unique instantiation of the traffic scenario, with differences arising from the randomized scenario configurations. These variations include different types of vehicles present in the scene, varying initial positions and trajectories of collaborative vehicles, difference in traffic flow and interactions, resulting in different visual, LiDAR inputs and control actions for each trail. Each trail contains RGB images and LiDAR point clouds from multiple vehicles, resulting in a substantially large dataset with a total size of approximately 400 GB.
@@ A.2.3 Behavior Cloning Details
In the CAV task, we apply Behavior Cloning (BC) following prior works COOPERNAUT [5] and STGN [8]. The process involves first running an expert agent to collect data from both the ego vehicle and collaborative vehicles. This data includes RGB images, LiDAR point clouds, and the control actions of the ego vehicle, specifically, braking decisions. Please see more details of the data collection in Appendix A.2.1.
@@ A.3.1 Experimental Setup
We resize the input RGB and depth images to $224\times 224$. To process RGB and depth data locally for each robot, we use ResNet-18 [14] as the encoder to extract feature maps of size $7\times 7$. The RGB features from both robots are shared and fused through channel-wise concatenation, and the depth features are processed similarly. Then we apply $1\times 1$ convolution to reduce the fused feature maps to the original channel dimensions for RGB and depth, respectively. We subsequently apply cross-attention to fuse the RGB and depth feature maps to generate multi-agent multi-modal feature aggregations. These aggregated features are passed through the decoder and upsampled to produce an output map matching the input image size.
@@ A.4 Training Complexity
We report the training complexity of AML [32] and CAML for the experiments of collaborative decision-making in CAV, and collaborative semantic segmentation for aerial-ground robots in Table 4 and Table 5, respectively. For the experiments, we employ a batch size of 32 and the Adam optimizer [18] with an initial learning rate of $1e{-3}$, and a Cosine Annealing Scheduler [24] to adjust the learning rate over time. The model is trained on an Nvidia RTX 3090 GPU with AMD Ryzen 9 5900 CPU and 32 GB RAM for 200 epochs.
@@ A.5 Knowledge Distillation
We begin by training a teacher decision-making model $\mathcal{T}$ offline using both RGB and LiDAR data, with a binary cross-entropy loss: $\mathcal{L}_{BCE}(y,\mathcal{T})=-\mathbb{E}_{\mathcal{D}}\big{[}y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\big{]}$, where $\mathcal{D}$ is the dataset, $y_{i}$ is the ground truth indicating whether the vehicle should brake, $p_{i}$ is the predicted probability by the teacher model $\mathcal{T}$. The student model $\mathcal{S}$ is trained to mimic the behavior of the teacher model while having less modalities. For each data point, the student model receives the same RGB image that the teacher model was given. The loss for the student model is a combination of two terms: the distillation loss using KL divergence between the student output and teacher output (soft targets), and the student task loss, which is the binary cross entropy loss between the student output and the true labels (hard targets). The soft targets from the teacher enrich learning with class similarities, while hard targets ensure alignment with true labels. The soft targets are generated by applying a temperature scaling to the logits. Following prior works [15, 32, 34], the scaled logits are defined as: $z_{i}=\frac{\exp{(z_{i}/t)}}{\exp{(z_{0}/t)}+\exp{(z_{1}/t)}}$, where $z_{i}$ is the logit for class $i$ and $t=4.0$ is the temperature. The distillation loss is defined as: $\mathcal{L}_{KD}(\mathcal{S},\mathcal{T})=-\sum z_{i}^{\mathcal{T}}\log(z_{i}^{\mathcal{S}})$, where $z_{i}^{\mathcal{T}}$ and $z_{i}^{\mathcal{S}}$ are the soft target probability from the teacher and student model, respectively. The overall loss for the student model is a weighted sum of the distillation loss and the binary cross-entropy loss: $\mathcal{L}_{\mathcal{S}}=(1-\alpha)\mathcal{L}_{BCE}(y,\mathcal{S})+\alpha t^{2}\mathcal{L}_{KD}(\mathcal{S},\mathcal{T})$ with $\alpha=0.9$ as the weight. After the training of knowledge distillation process, we obtain a student model that uses only RGB data while learning from a teacher model that has access to both RGB and LiDAR data. This enables the student model to be effective during testing with only RGB data. Additionally, by leveraging knowledge distillation, the student model benefits from the additional insights provided by the LiDAR data during training, learning more effectively compared to training solely with RGB data.
@@ A.6 Limitations and Future Work
CAML can handle different and mixed modalities for the agents at test time. Feature embeddings are first extracted by modality-specific encoders and then fused across agents (e.g., via concatenation or cross-attention) for downstream decision-making. In the student model of CAML, the modalities for each agent can be configured individually. The agents do not all need to have the same modalities. The current student model training does not dynamically change the modalities over time. However, we can handle dynamic changes without changing the architecture. Our planned solution is to incorporate a modality dropout strategy. During training of knowledge distillation, we randomly drop out or mask some modality for each agent in the student model. In this way, the student model learns to handle cases where some modalities may vary dynamically for different agents. Exploring this strategy would be a promising direction for future work. Also similar to prior works (e.g., STGN [8], Who2com [22], When2com [21], and Where2comm [17]), CAML assumes reliable communication during training. While this is common in the field, future work could explore robustness under stochastic communication delay or failure.

### v3

@@ 4.1 Collaborative Decision-Making
[…]
Next, we evaluate the case involving more modalities beyond RGB and LiDAR. Each vehicle now receives RGB, LiDAR, and state information (e.g., position and velocity) during training, while only RGB inputs are available during inference. As shown in Table 3, incorporating state information as an additional modality further improves system performance, demonstrating CAML’s ability to effectively leverage richer multimodal data during training while maintaining strong performance under reduced-modality conditions.
Table 3: Performance evaluation with three modalities: RGB, LiDAR, and state information. Incorporating state information as an additional modality further enhances overall system performance.
Modalities Overtaking Left Turn Red Light Violation
[…]
We also investigate the complementary role of different modalities by retaining LiDAR and removing RGB during testing. As presented in Table 4, the performance of retaining LiDAR is slightly better than retaining RGB, as LiDAR provides more precise spatial localization, which is beneficial for accident detection.
Table 4: Performance comparison (%) of CAML when retaining either RGB or LiDAR during testing. Both settings achieve comparable performance, but retaining LiDAR yields slightly better results due to its more precise spatial localization, which benefits accident detection.
Modality Overtaking Left Turn Red Light Violation
[…]
Furthermore, we compare CAML, which adopts an intermediate cooperation strategy, with a late cooperation baseline. In the late cooperation setting, each vehicle independently processes RGB and LiDAR inputs using the same encoders as in CAML (ResNet-18 for RGB and PointTransformer for LiDAR), followed by MLP heads to generate action logits. For each modality, we first fuse the logits from connected vehicles via averaging and then fuse the averaged logits across modalities to determine the ego vehicle’s control actions. We apply the same KD procedure as in CAML, distilling from the RGB–LiDAR teacher model to a student model that uses only RGB at test time. As shown in Table 5, CAML outperforms the late cooperation approach by enabling the model to learn richer and more complementary representations across agents and modalities, whereas late cooperation relies solely on output-level fusion, leading to potential information loss.
Table 5: Performance comparison (%) between CAML and the late cooperation approach. CAML achieves superior results by enabling the model to learn richer and more complementary representations across agents and modalities.
Approach Overtaking Left Turn Red Light Violation
[…]
@@ 4.2 Collaborative Semantic Segmentation
To further evaluate our approach, we focus on collaborative semantic segmentation by conducting experiments with real-world data from aerial-ground robots. We use the dataset CoPeD [zhou2024coped], with one aerial robot and one ground robot, in two different real-world scenarios of the indoor NYUARPL and the outdoor HOUSEA. For more details about the dataset, please refer to [zhou2024coped]. Additionally, we introduce noise to the RGBD data collected by the ground robot. For both aerial and ground robots, RGB and depth data are used during training, while only RGB data is used during testing in CAML.
@@
We adopt the FCN [long2015fully] architecture as the backbone for semantic segmentation. To process RGB and depth data locally for each robot, we use ResNet-18 [he2016deep] as the encoder to extract feature maps of size $7\times 7$. Please refer to Appendix A.3.1 for more details of the experimental setup. We first train a teacher model offline with aerial-ground robots collaboration using cross-entropy loss, where each robot has both RGB and depth data. Then we train a student model to mimic the behavior of the teacher model with only RGB data for both aerial and ground robots through KD. The KD process is similar to that of the collaborative decision-making in CAV, but here we use a cross-entropy loss as the student task loss. Please see Appendix A.4 for the detailed training settings. We evaluate performance using the Mean Intersection over Union (mIoU) metric, which quantifies the average overlap between predicted segmentation outputs and ground truth across all classes. We compare the performance of CAML with other baselines including AML [shen2023auxiliary] and FCN [long2015fully]. In the AML approach, only the ground robot operates, with RGB and depth data available during training but only RGB data used for testing. The FCN approach involves only the ground robot operating with RGB data for both training and testing.
Table 6: Baseline Comparison of Semantic Segmentation on real-world dataset CoPeD [zhou2024coped] using aerial-ground robots in indoor and outdoor environments. CAML achieves the highest mIoU in both environments, with upto 10.6% higher accuracy.
@@
Experimental Results.
We first present the experimental results of baselines comparison in Table 6, where CAML demonstrates superior performance in terms of mIoU across both indoor and outdoor environments. Specifically, CAML achieves an improvement of mIoU for $\bf 8.9\%$} in indoor scenario and $\bf 10.6\%$} in outdoor scenario compared to AML [shen2023auxiliary]. We also show the qualitative results in Fig. 5. As we can see, despite the noisy input image from the ground robot, CAML produces predictions that are closest to the ground truth. This improvement is attributed to CAML’s multi-agent collaboration, which provides complementary information to enhance data coverage and offers a more comprehensive understanding of the scenes. Additionally, the utilization of auxiliary depth data during training results in more precise segmentation outputs.
@@ 5 Conclusions
In conclusion, we propose Collaborative Auxiliary Modality Learning (CAML), a unified framework for multi-modal multi-agent systems. Unlike prior methods that either focus on multi-agent collaboration without modality reduction or address multi-modal learning in single-agent settings, CAML integrates both aspects. It enables agents to collaborate using shared modalities during training while allowing efficient, modality-reduced inference. This not only lowers computational costs and data requirements at test time but also enhances predictive accuracy through multi-agent collaboration. We provide an intuitive analysis of CAML in terms of data coverage and information gain, justifying its effectiveness. CAML demonstrates up to a $58.1\%$ improvement in accident detection for connected autonomous driving in complex scenarios and up to a $10.6\%$ mIoU gain in real-world aerial-ground collaborative semantic segmentation. These improvements underscore the practical implications of our framework, especially for resource-constrained environments. For a discussion on limitations and future work, please see Appendix A.6.
Acknowledgment
This research is supported in part by National Science Foundation, Dr. Barry Mersky & Capital One E-Nnovate Endowed Professorships.
@@ A.2.1 Dataset Details
The dataset is generated using the AUTOCASTSIM [qiu2021autocast] benchmark, which features three complex and accident-prone traffic scenarios for connected autonomous driving, characterized by limited sensor coverage or obstructed views. These scenarios are realistic and include background traffic of 30 vehicles. They involve challenging interactions such as overtaking, lane changing, and red-light violations, which inherently increase the risk of accidents: (1) Overtaking: A sedan is blocked by a truck on a narrow, two-way road with a dashed centerline. The truck also obscures the sedan’s view of oncoming traffic. The ego vehicle must decide when and how to safely pass the truck. (2) Left Turn: The ego vehicle attempts a left turn at a yield sign. Its view is partially blocked by a truck waiting in the opposite left-turn lane, reducing visibility of vehicles coming from the opposite direction. (3) Red Light Violation: As the ego vehicle crosses an intersection, another vehicle runs a red light. Due to nearby vehicles waiting to turn left, the ego vehicle’s sensors are unable to detect the violator.
Following the setup established by prior works COOPERNAUT [cui2022coopernaut] and STGN [gao2024collaborative], we collect 24 data trails for each scenario, using 12 trails for training and the remaining 12 for testing. Each trail represents a unique instantiation of the traffic scenario, with differences arising from the randomized scenario configurations. These variations include different types of vehicles present in the scene, varying initial positions and trajectories of collaborative vehicles, difference in traffic flow and interactions, resulting in different visual, LiDAR inputs and control actions for each trail. Each trail contains RGB images and LiDAR point clouds from multiple vehicles, resulting in a substantially large dataset with a total size of approximately 400 GB.
@@ A.2.3 Behavior Cloning Details
In the CAV task, we apply Behavior Cloning (BC) following prior works COOPERNAUT [cui2022coopernaut] and STGN [gao2024collaborative]. The process involves first running an expert agent to collect data from both the ego vehicle and collaborative vehicles. This data includes RGB images, LiDAR point clouds, and the control actions of the ego vehicle, specifically, braking decisions. Please see more details of the data collection in Appendix A.2.1.
@@ A.3.1 Experimental Setup
We resize the input RGB and depth images to $224\times 224$. To process RGB and depth data locally for each robot, we use ResNet-18 [he2016deep] as the encoder to extract feature maps of size $7\times 7$. The RGB features from both robots are shared and fused through channel-wise concatenation, and the depth features are processed similarly. Then we apply $1\times 1$ convolution to reduce the fused feature maps to the original channel dimensions for RGB and depth, respectively. We subsequently apply cross-attention to fuse the RGB and depth feature maps to generate multi-agent multi-modal feature aggregations. These aggregated features are passed through the decoder and upsampled to produce an output map matching the input image size.
@@ A.4 Training Complexity
We report the training complexity of AML [shen2023auxiliary] and CAML for the experiments of collaborative decision-making in CAV, and collaborative semantic segmentation for aerial-ground robots in Table 7 and Table 8, respectively. For the experiments, we employ a batch size of 32 and the Adam optimizer [kingma2014adam] with an initial learning rate of 1e31e{-3}, and a Cosine Annealing Scheduler [loshchilov2016sgdr] to adjust the learning rate over time. The model is trained on an Nvidia RTX 3090 GPU with AMD Ryzen 9 5900 CPU and 32 GB RAM for 200 epochs.
@@ A.5 Knowledge Distillation
We begin by training a teacher decision-making model $\mathcal{T}$ offline using both RGB and LiDAR data, with a binary cross-entropy loss: $\mathcal{L}_{BCE}(y,\mathcal{T})=-\mathbb{E}_{\mathcal{D}}\big[y_{i}\log(p_{i})+(1-y_{i})\log(1-p_{i})\big]$, where $\mathcal{D}$ is the dataset, $y_{i}$ is the ground truth indicating whether the vehicle should brake, $p_{i}$ is the predicted probability by the teacher model $\mathcal{T}$. The student model $\mathcal{S}$ is trained to mimic the behavior of the teacher model while having less modalities. For each data point, the student model receives the same RGB image that the teacher model was given. The loss for the student model is a combination of two terms: the distillation loss using KL divergence between the student output and teacher output (soft targets), and the student task loss, which is the binary cross entropy loss between the student output and the true labels (hard targets). The soft targets from the teacher enrich learning with class similarities, while hard targets ensure alignment with true labels. The soft targets are generated by applying a temperature scaling to the logits. Following prior works [hinton2015distilling, shen2023auxiliary, tian2019contrastive], the scaled logits are defined as: $z_{i}=\frac{\exp{(z_{i}/t)}}{\exp{(z_{0}/t)}+\exp{(z_{1}/t)}}$, where $z_{i}$ is the logit for class $i$ and $t=4.0$ is the temperature. The distillation loss is defined as: $\mathcal{L}_{KD}(\mathcal{S},\mathcal{T})=-\sum z_{i}^{\mathcal{T}}\log(z_{i}^{\mathcal{S}})$, where $z_{i}^{\mathcal{T}}$ and $z_{i}^{\mathcal{S}}$ are the soft target probability from the teacher and student model, respectively. The overall loss for the student model is a weighted sum of the distillation loss and the binary cross-entropy loss: $\mathcal{L}_{\mathcal{S}}=(1-\alpha)\mathcal{L}_{BCE}(y,\mathcal{S})+\alpha t^{2}\mathcal{L}_{KD}(\mathcal{S},\mathcal{T})$ with $\alpha=0.9$ as the weight. After the training of knowledge distillation process, we obtain a student model that uses only RGB data while learning from a teacher model that has access to both RGB and LiDAR data. This enables the student model to be effective during testing with only RGB data. Additionally, by leveraging knowledge distillation, the student model benefits from the additional insights provided by the LiDAR data during training, learning more effectively compared to training solely with RGB data.
@@ A.6 Limitations and Future Work
CAML is capable of handling different and mixed modalities across agents at test time. Feature embeddings are first extracted using modality-specific encoders and then fused across agents (e.g., via concatenation or cross-attention) for downstream decision-making. In the student model of CAML, the modality configuration can be specified individually for each agent, allowing flexibility without requiring all agents to share the same modalities. Although the current student model does not dynamically vary modalities over time, such capability can be achieved without modifying the architecture. A promising direction is to incorporate a modality dropout strategy during knowledge distillation, where certain modalities are randomly masked or removed for each agent during training. This would enable the student model to better adapt to dynamically changing modality availability. Moreover, CAML can be extended with uncertainty estimation [liu2023data, liu2024towards] to assess the reliability of each modality and enhance robustness when sensor inputs are noisy or degraded (e.g., under adverse weather conditions in connected autonomous driving).